2C4H10(g)+13O2(g)→8CO2(g)+10H2O(g)   ∆H= -2877 kJ

How much energy is produced when 723.4 g of camping gas is burnt? 


1
Expert's answer
2021-04-01T02:27:56-0400

Solution.

2C4H10(g)+13O2(g)→8CO2(g)+10H2O(g) ∆H= -2877 kJ


First, we calculate the amount of substance for the gas. We will do this using this formula:

"n=m\/M" (1)

(where "n" is the amount of substance, "m" is its mass, "M" is the molar mass)


We find the molar mass of the gas using the atomic masses from the Periodic Table of Elements:

"M=Ar(C)*4+Ar(H)*10" (2)

"M=4*12+1*10=58" g/mol (3)

"n=723.4\/58=12.47" mol (4)


When ∆H is equal to a negative number, this means that energy is realized into the environment +Q. Now we can calculate how much energy was realized using the proportion:

"2 C4H10\/12.47 C4H10=2877kJ\/x" (5)

Where does x equal:

"x=12.47*2877\/2=17 938 kJ"


Answer: Q=17938 kJ